# Hebbian Learning from First Principles

### v1

@@ 1 Introduction
Since Parisi’s groundbreaking discovery of multiple equilibria in complex systems GiorgioOld ; GiorgioNobel , spin glasses (i.e. complex systems tackled by statistical mechanical techniques MPV ; FisherHertz ) were immediately recognized as suitable systems for modelling the spontaneous information processing capabilities collectively shown by assemblies of neurons in neural network models Coolen ; Nishimori ; Haiping ; Engel ; Kadmon ; LenkaJPA ; storage2 . The first celebrated statistical mechanical analysis of a spin glass model for neural networks was addressed in the eighties by Amit, Gutfreund and Sompolinsky (resulting in the so-called AGS theory) AGS ; Amit , focusing on the Hopfield model Hopfield . The latter is a fully connected mean field neural network whose neurons interact in couples through a synaptic coupling that implements the celebrated Hebbian prescription neurons that fire together wire together.
@@
Next, by relying upon Guerra’s interpolation GuerraNN ; AABO-JPA2020 ; guerra_broken ; Jean0 , we prove that, in the limit of infinite information provided to the network (a scenario that we call big data limit super ; unsup ), whatever the learning procedure (i.e., with or without the presence of a teacher) these Hebbian learning rules converge to the original Hopfield prescription for pattern storage [^2] Strictly speaking, we have convergence to the Hopfield network solely if neural interactions are pairwise, while the dense counterpart converges to the Hebbian P-spin-glass (studied by Elisabeth Gardner in Statistical Mechanics Gardner and Pierre Baldi in Computer Science Baldi ), as it should be.: we prove this both at the level of the Cost functions as well as involving the whole free energy. Thus, in this limit, AGS theory AGS is entirely recovered. Finally, we also prove a one-to-one relation between these new Hamiltonians (i.e., Cost functions), usually preferred in Statistical Mechanics and standard Loss functions, typically used in Machine Learning, so to extend the results obtained for the former to the latter.
@@ 2.1 The Hopfield neural network in statistical mechanics
(Control and order parameters) Given the Hopfield model, we denote by $\beta\in\mathbb{R}^{+}$ the noise in the network (see below in eq. (6)) and by $\alpha:=K/N\in\mathbb{R}^{+}$ the storage capacity of the network.
@@
$m_{\mu}$ $=\dfrac{1}{N}\sum_{i=1}^{N}\xi^{\mu}_{i}\sigma_{i},$ (4)
@@
Roughly speaking, control parameters are parameters that we freely tune in the network (and we expect that -as for the former- the higher the noise in the network, the harder the computation for its neurons -whatever it is, storage or retrieval- much as for the latter, because the larger the storage, the harder the task). At contrary, order parameters are response functions, that is, they provide a quantitative measure of network’s performances as control parameters vary.
@@
$H_{N}^{\textnormal{(Hop)}}(\bm{\sigma}|J_{0}=1,h=0,\bm{\xi})=-\frac{N}{2}\sum_{\mu=1}^{K}m_{\mu}^{2}.$ (5)
@@
$\lim_{t\to\infty}\mathbb{P}(\bm{\sigma}(t)|\bm{\xi})=\mathbb{P}(\bm{\sigma}|\bm{\xi})=\frac{e^{-\beta H_{N}^{\textnormal{(Hop)}}(\bm{\sigma}|\bm{\xi})}}{\sum_{\bm{\sigma}}e^{-\beta H_{N}^{\textnormal{(Hop)}}(\bm{\sigma}|\bm{\xi})}}=\frac{e^{\frac{N\beta}{2}\sum_{\mu}^{K}m_{\mu}^{2}}}{Z_{N}(\bm{\xi})}$ (8)
@@
It is rather easy to understand at least the order of magnitude of the Hopfield storage, i.e. $K\propto N^{1}$, by a signal-to-noise argument that we now streamline: as we want the patterns to be configurations that are stable points, i.e. the attractors for the neural dynamics provided in eq. (6), let us assume that the network already lies in a pattern and checks for its stability. We set now $\bm{\sigma}(t)=\bm{\xi}^{1}$ and $h=0$ (no bias must be added): if $\bm{\xi}^{1}$ is an attractor, it must happen that $\sigma_{i}(t+1)=\sigma_{i}(t)$ for all $i=1,...,N$. Since the stochastic process described in eq. (6) prescribes $\sigma_{i}(t+1)=sign[\tanh(\beta\sum_{j}J_{ij}\sigma_{j}(t))]$, by substituting in this formula the expression the coupling $J_{ij}$ provided in eq. (3) and the assumption $\sigma_{j}(t)=\xi_{j}^{1}$, we have that $\sigma_{i}(t+1)=sign[\tanh(\beta N^{-1}\sum_{j=1}^{N}\sum_{\mu=1}^{K}\xi_{i}^{\mu}\xi_{j}^{\mu}\xi_{j}^{\mbox{\tiny$1$}}]$: this sign is governed by $N^{-1}\sum_{\mu=1}^{K}\xi_{i}^{\mu}\xi_{j}^{\mu}\xi_{j}^{\mbox{\tiny$1$}}\sim\xi_{i}^{1}\pm\sqrt{K/N}=\xi_{i}^{1}\pm\sqrt{\alpha}$, where the contribution $\xi_{i}^{1}$ plays as the signal while the storage plays as a noise. Thus, as long as the storage $\alpha<1$, the value of $\sigma_{i}(t+1)$ is the sign of $\xi_{i}^{1}$, that, in turn, is also the value for $\sigma_{i}(t)$: hence patterns can play as attractors for neural dynamics provided they are not too many ($\alpha<1$). Note that, from an Information Theory perspective, when the network reaches a stationary state (i.e., an attractor representing a stored pattern), the final state of the stochastic variable $\sigma_{i}$ directly codes for the $i^{th}$ bit of information of the corresponding pattern that it is retrieving (e.g., $\sigma_{i}=\xi_{i}^{1}$ in the example above).
The argument of Remark 1 however is rather rough and, in order to have a sharp estimate about the value of the maximal storage $\alpha_{c}\approx 0.14$, one has to solve for the free energy (that we just define hereafter while we remind to standard textbook for its explicit evaluation Amit ; Coolen ) and inspect its singularities (that depict computational phase transitions and thus automatically detect the maximal storage $\alpha_{c}$). The free energy, defined as the intensive logarithm of the partition function, is the main observable in Statistical Mechanics and its extremization accounts for the balance between energy minimization and entropy maximization (at given noise $\beta$ and storage $\alpha$), or -in Machine Learning jargon- it allows to search for the minima of the Cost function under prescription of the maximum entropy, framed as an inferential criterion.
@@ 2.2 The Hopfield neural network from statistical inference
The brackets represent the Boltzmann average over the ensemble $\omega(.)$ and over the quenched patterns $\xi$, namely -given a generic observable $O(\sigma,\xi)$ (function of both the neural activities and the patterns lying in the synaptic coupling)- we write
@@ 3 Main theme: Hebbian learning
As discussed in the Introduction, despite the expression of the coupling provided in eq. (3) is often named Hebbian learning, the above model has little to share with modern Machine Learning as there is no real learning process underlying the storage of patterns within the synaptic matrix of the Hopfield neural network. At contrary, mimicking the Machine Learning literature, we would only provide the network with examples of the patterns (and never the patterns themselves) in order to check if and how, once enough examples have been supplied, the network is able to reconstruct the patterns hidden in the information it experienced. To do so, we need also to generalize the synaptic coupling in the r.h.s. of eq. (2) in order to let the network experience examples of patterns, rather than pattern themselves. Let us first generate such a dataset of examples.
@@
(Control and order parameters) Beyond the noise in the network $\beta\in\mathcal{R}^{+}$ and the storage capacity of the network $\alpha:=K/N\in\mathcal{R}^{+}$ (as in the previous model at work solely with pattern storage), further control parameters for Hebbian learning are also the dataset quality $r$ and the dataset quantity $M$: note that these two variables merge together to form $\rho=(1-r^{2})/Mr^{2}$ that is the entropy of the dataset [^6] Stricktly speaking, $\rho:=(1-r^{2})/Mr^{2}$ is the conditional entropy of the pattern given the examples, i.e., $S(\xi_{i}^{\mu}|\bm{\eta_{i}^{\mu}})$. The latter accounts for the amount of information required to describe the pattern entry $\xi_{i}^{\mu}$ given the knowledge of $M$ independent observations of (examples of) the pattern $\bm{\eta}_{i}^{\mu}$ $=(\eta_{i}^{\mu,a=1},...,\eta_{i}^{\mu,a=M})$. While we deepened this point elsewhere (see e.g. prlmiriam ), to see that $\rho$ quantifies the conditional entropy of the pattern given the examples $S(\xi_{i}^{\mu}|\bm{\eta_{i}^{\mu}})$, an intuitive argument is to note that the error probability per bit is $\mathbb{P}(\chi_{i}^{\mu,a}=-1)=(1-r)/2$, thus by applying the majority rule to the empirical vector $\bm{\eta}_{i}^{\mu}$ we get $\mathbb{P}(sign(\sum_{a=1}^{M}\chi_{i}^{\mu,a}=-1)\sim 1-erf(1/\sqrt{2\rho})$. Further, it is simple to check that the dataset is maximally informative when its entropy is null, namely $\rho=0$, that is or $r=1$ (i.e. examples are perfect copies of the patterns) or $M\to\infty$ (i.e. examples are infinite)..
@@
It is a trivial exercise to prove that the Cost functions provided in Def. 6 and Def. 7 can indeed be written as quadratic forms, respectively, in the order parameters provided by eq. (21) and eq. (22). Further, it is elementary to realize that -when moving from storing to learning- other questions beyond the maximal storage capacity of the network can be addressed, at first the thresholds for learning, namely the minimal amount of information that must be provided to the network in order for it to correctly infer the patterns: these computational aspects are known and they have been discussed elsewhere prlmiriam ; Lucibello .
@@ 3.1 Maximum entropy principle for shallow neural networks
The MEP Jaynes allows to determine the less structured probability distribution whose moments best fit their empirical counterparts [^7] Note that the empirical moments have to be experimentally measured as for instance discussed for the derivation of the Hopfield storage prescription, see Sec. 2.2. See also, e.g. Application1 ; Application2 ; Application3 for concrete applications of this method to real datasets collected in biological sciences (neuroscience, ecology and immunology respectively).. As we are going to show in the first part of this manuscript, as long as we apply this technique to pairwise networks (as the Hopfield model where neurons interact in couples), these moments turn out to be just the first-order moments (i.e. means and variances or one-point and two-point correlation functions). At contrary, in the second part of this manuscript, we show how, extending the theory to many body interactions, higher order moments have to be accounted.
@@ 3.2.1 Supervised Hebbian learning
We apply the CLT to the random variable $\dfrac{1}{M}\displaystyle\sum\limits_{a=1}^{M}\eta_{i}^{\mu,a}$ in the $M\to+\infty$ limit. With some trivial manipulations we can state that
@@
Focusing on $\mathbb{P}_{N,M}(\bm{\sigma}|\bm{\eta})$ in eq. (25) and, in particular, by applying the Central Limit Theorem to the random variable $\dfrac{1}{M}\displaystyle\sum\limits_{a=1}^{M}\eta_{i}^{\mu,a}$ in the $M\to+\infty$ limit, as already done in Lemma 1, its probability distribution becomes a Dirac’s delta peaked in $\xi^{\mu}_{i}$ and, thus, we reach the thesis. ∎
@@ 3.4 Maximum entropy principle for dense neural networks
Roughly speaking, it is reasonable to expect that -as long as the neurons interact in couples (see, e.g. (2))- the Hamiltonians related to the networks are quadratic forms in the order parameters (the $m_{\mu}$ for the Hopfield case, see Def. 8), hence the Boltzmann-Gibbs probabilities that the MEP prescribes to them have Gaussian shape (see e.g. eq. (8)) and, consistently, solely means and variances in the datasets contribute to their formation (see e.g. eq. (13)).

### v2

@@ 1 Introduction
Since Parisi’s groundbreaking discovery of multiple equilibria in complex systems GiorgioOld ; GiorgioNobel , spin glasses (i.e. complex systems tackled by statistical mechanical techniques MPV ; FisherHertz ) were immediately recognized as suitable systems for modelling the spontaneous information processing capabilities collectively shown by assemblies of neurons in neural network models Coolen ; Nishimori ; Haiping ; Engel ; Kadmon ; LenkaJPA ; storage2 . The first celebrated statistical mechanical analysis of a spin glass model for neural networks was addressed in the eighties by Amit, Gutfreund and Sompolinsky (resulting in the so-called AGS theory) AGS ; Amit , focusing on the Hopfield model Hopfield . The latter is a fully connected mean field neural network whose neurons interact in couples through a synaptic coupling that implements the celebrated Hebbian prescription neurons that fire together wire together Hebb .
@@
Next, by relying upon Guerra’s interpolation GuerraNN ; guerra_broken , we prove that, in the limit of infinite information provided to the network (a scenario that we call big data limit super ; unsup ), whatever the learning procedure (i.e., with or without the presence of a teacher) these Hebbian learning rules converge to the original Hopfield prescription for pattern storage [^2] Strictly speaking, we have convergence to the Hopfield network solely if neural interactions are pairwise, while the dense counterpart converges to the Hebbian P-spin-glass (studied by Elisabeth Gardner in Statistical Mechanics Gardner and Pierre Baldi in Computer Science Baldi ), as it should be.. We prove this both at the level of the Cost functions as well as involving the whole free energy. Thus, in this limit, AGS theory AGS is entirely recovered. Guerra’s interpolation is a mathematical technique introduced for the first time by Francesco Guerra guerra2002thermodynamic for spin glasses and later imported to neural networks AABO-JPA2020 ; Jean0 . It is mathematically justified and, in general, easier than the so-known replica trick MPV . Finally, we also prove a one-to-one relation between these new Hamiltonians (i.e., Cost functions), usually preferred in Statistical Mechanics and standard Loss functions, typically used in Machine Learning, so to extend the results obtained for the former to the latter.
@@ 2.1 The Hopfield neural network in statistical mechanics
(Control and order parameters) Given the Hopfield model, the control parameters are $\beta\in\mathbb{R}^{+}$, the noise (see below in eq. (6)), and $\alpha:=K/N\in\mathbb{R}^{+}$, the storage capacity of the network.
@@
$m_{\mu}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\bm{\sigma}|\bm{\xi})}$ $=\dfrac{1}{N}\sum_{i=1}^{N}\xi^{\mu}_{i}\sigma_{i},$ (4)
@@
Roughly speaking, control parameters are parameters that we freely tune in the network (and we expect that –as for the former– the higher the noise in the network, the harder the computation for its neurons –whatever it is, storage or retrieval– much as for the latter, because the larger the storage, the harder the task). On the contrary, order parameters are response functions, that is, they provide a quantitative measure of network’s performances as control parameters vary.
@@
$H_{N}^{\textnormal{(Hop)}}(\bm{\sigma}|J_{0}=1,h=0,\bm{\xi})=-\frac{N}{2}\sum_{\mu=1}^{K}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(m_{\mu}(\bm{\sigma}|\bm{\xi}))^{2}}.$ (5)
@@
$\lim_{t\to\infty}\mathbb{P}(\bm{\sigma}(t)|\bm{\xi})=\mathbb{P}(\bm{\sigma}|\bm{\xi})=\frac{e^{-\beta H_{N}^{\textnormal{(Hop)}}(\bm{\sigma}|\bm{\xi})}}{\sum_{\bm{\sigma}}e^{-\beta H_{N}^{\textnormal{(Hop)}}(\bm{\sigma}|\bm{\xi})}}=\frac{e^{\frac{N\beta}{2}\sum_{\mu}^{K}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(m_{\mu}(\bm{\sigma}|\bm{\xi}))^{2}}}}{Z_{N}(\bm{\xi})}$ (8)
@@
It is rather easy to understand at least the order of magnitude of the Hopfield storage, i.e. $K\propto N^{1}$, by a signal-to-noise argument that we now streamline: as we want the patterns to be configurations that are stable points, i.e. the attractors for the neural dynamics provided in eq. (6), let us assume that the network already lies in a pattern and checks for its stability. We set now $\bm{\sigma}(t)=\bm{\xi}^{1}$ and $h=0$ (no bias must be added): if $\bm{\xi}^{1}$ is an attractor, it must happen that $\sigma_{i}(t+1)=\sigma_{i}(t)$ for all $i=1,...,N$. Since the stochastic process described in eq. (6) prescribes $\sigma_{i}(t+1)=\textrm{sign}[\tanh(\beta\sum_{j}J_{ij}\sigma_{j}(t))]$, by substituting in this formula the expression the coupling $J_{ij}$ provided in eq. (3) and the assumption $\sigma_{j}(t)=\xi_{j}^{1}$, we have that $\sigma_{i}(t+1)=\textrm{sign}[\tanh(\beta N^{-1}\sum_{j=1}^{N}\sum_{\mu=1}^{K}\xi_{i}^{\mu}\xi_{j}^{\mu}\xi_{j}^{\mbox{\tiny$1$}}]$: this sign is governed by $N^{-1}\sum_{\mu=1}^{K}\xi_{i}^{\mu}\xi_{j}^{\mu}\xi_{j}^{\mbox{\tiny$1$}}\sim\xi_{i}^{1}\pm\sqrt{K/N}=\xi_{i}^{1}\pm\sqrt{\alpha}$, where the contribution $\xi_{i}^{1}$ plays as the signal while the storage plays as a noise. Thus, as long as the storage $\alpha<1$, the value of $\sigma_{i}(t+1)$ is the sign of $\xi_{i}^{1}$, that, in turn, is also the value for $\sigma_{i}(t)$: hence patterns can play as attractors for neural dynamics provided they are not too many ($\alpha<1$). Note that, from an Information Theory perspective, when the network reaches a stationary state (i.e., an attractor representing a stored pattern), the final state of the stochastic variable $\sigma_{i}$ directly codes for the $i^{th}$ bit of information of the corresponding pattern that it is retrieving (e.g., $\sigma_{i}=\xi_{i}^{1}$ in the example above).
The argument of Remark 1 however is rather rough and, in order to have a sharp estimate about the value of the maximal storage $\alpha_{c}\approx 0.14$, one has to solve for the free energy (that we just define hereafter while we remind to standard textbook for its explicit evaluation Amit ; Coolen ) and inspect its singularities (that depict computational phase transitions and thus automatically detect the maximal storage $\alpha_{c}$). The free energy, defined as the intensive logarithm of the partition function, is the main observable in Statistical Mechanics and its extremization accounts for the balance between energy minimization and entropy maximization (at given noise $\beta$ and storage $\alpha$), or –in Machine Learning jargon– it allows to search for the minima of the Cost function under prescription of the maximum entropy, framed as an inferential criterion.
@@ 2.2 The Hopfield neural network from statistical inference
The brackets represent the Boltzmann average over the ensemble $\omega(.)$ and over the quenched patterns $\xi$, namely –given a generic observable $O(\sigma,\xi)$ (function of both the neural activities and the patterns lying in the synaptic coupling)– we write
@@ 3 Main theme: Hebbian learning
As discussed in the Introduction, despite the expression of the coupling provided in eq. (3) is often named Hebbian learning, the above model has little to share with modern Machine Learning as there is no real learning process underlying the storage of patterns within the synaptic matrix of the Hopfield neural network. On the contrary, mimicking the Machine Learning literature, we would only provide the network with examples of the patterns (and never the patterns themselves) in order to check if and how, once enough examples have been supplied, the network is able to reconstruct the patterns hidden in the information it experienced. To do so, we need also to generalize the synaptic coupling in the r.h.s. of eq. (2) in order to let the network experience examples of patterns, rather than pattern themselves. Let us first generate such a dataset of examples.
@@
(Control and order parameters) Beyond the noise in the network $\beta\in\mathcal{R}^{+}$ and the storage capacity of the network $\alpha:=K/N\in\mathcal{R}^{+}$ (as in the previous model at work solely with pattern storage), further control parameters for Hebbian learning are also the dataset quality $r$ and the dataset quantity $M$: note that these two variables merge together to form $\rho=(1-r^{2})/Mr^{2}$ that is the entropy of the dataset [^6] Stricktly speaking, $\rho:=(1-r^{2})/Mr^{2}$ is the conditional entropy of the pattern given the examples, i.e., $S(\xi_{i}^{\mu}|\bm{\eta_{i}^{\mu}})$. The latter accounts for the amount of information required to describe the pattern entry $\xi_{i}^{\mu}$ given the knowledge of $M$ independent observations of (examples of) the pattern $\bm{\eta}_{i}^{\mu}$ $=(\eta_{i}^{\mu,a=1},...,\eta_{i}^{\mu,a=M})$. While we deepened this point elsewhere (see e.g. prlmiriam ), to see that $\rho$ quantifies the conditional entropy of the pattern given the examples $S(\xi_{i}^{\mu}|\bm{\eta_{i}^{\mu}})$, an intuitive argument is to note that the error probability per bit is $\mathbb{P}(\chi_{i}^{\mu,a}=-1)=(1-r)/2$, thus by applying the majority rule to the empirical vector $\bm{\eta}_{i}^{\mu}$ we get $\mathbb{P}(\textrm{sign}(\sum_{a=1}^{M}\chi_{i}^{\mu,a}=-1)\sim 1-erf(1/\sqrt{2\rho})$. Further, it is simple to check that the dataset is maximally informative when its entropy is null, namely $\rho=0$, that is or $r=1$ (i.e. examples are perfect copies of the patterns) or $M\to\infty$ (i.e. examples are infinite)..
@@
It is a trivial exercise to prove that the Cost functions provided in Def. 6 and Def. 7 can indeed be written as quadratic forms, respectively, in the order parameters provided by eq. (21) and eq. (22). Further, it is elementary to realize that –when moving from storing to learning– other questions beyond the maximal storage capacity of the network can be addressed, at first the thresholds for learning, namely the minimal amount of information that must be provided to the network in order for it to correctly infer the patterns: these computational aspects are known and they have been discussed elsewhere prlmiriam ; Lucibello .
@@ 3.1 Maximum entropy principle for shallow neural networks
The MEP Jaynes allows to determine the less structured probability distribution whose moments best fit their empirical counterparts [^7] Note that the empirical moments have to be experimentally measured as for instance discussed for the derivation of the Hopfield storage prescription, see Sec. 2.2. See also, e.g. Application1 ; Application2 ; Application3 for concrete applications of this method to real datasets collected in biological sciences (neuroscience, ecology and immunology respectively).. As we are going to show in the first part of this manuscript, as long as we apply this technique to pairwise networks (as the Hopfield model where neurons interact in couples), these moments turn out to be just the first-order moments (i.e. means and variances or one-point and two-point correlation functions). On the contrary, in the second part of this manuscript, we show how, extending the theory to many body interactions, higher order moments have to be accounted.
@@ 3.1.1 Supervised Hebbian learning
The independence of $\lambda_{2}$ by the index $\mu$ reflects that the weights allocated to each pattern (i.e. the amplitudes of their basins of attraction) are all the same.
@@ 3.2.1 Supervised Hebbian learning
We apply the Central Limit Theorem (CLT) to the random variable $\dfrac{1}{M}\displaystyle\sum\limits_{a=1}^{M}\eta_{i}^{\mu,a}$ in the $M\to+\infty$ limit. With some trivial manipulations we can state that
@@
Focusing on $\mathbb{P}_{N,M}(\bm{\sigma}|\bm{\eta})$ in eq. (25) and, in particular, by applying the CLT to the random variable $\dfrac{1}{M}\displaystyle\sum\limits_{a=1}^{M}\eta_{i}^{\mu,a}$ in the $M\to+\infty$ limit, as already done in Lemma 1, its probability distribution becomes a Dirac’s delta peaked in $\xi^{\mu}_{i}$ and, thus, we reach the thesis. ∎
@@ 3.4 Maximum entropy principle for dense neural networks
Roughly speaking, it is reasonable to expect that –as long as the neurons interact in couples (see, e.g. (2))– the Hamiltonians related to the networks are quadratic forms in the order parameters (the $m_{\mu}$ for the Hopfield case, see Def. 8), hence the Boltzmann-Gibbs probabilities that the MEP prescribes to them have Gaussian shape (see e.g. eq. (8)) and, consistently, solely means and variances in the datasets contribute to their formation (see e.g. eq. (13)).
@@
Note that, while these dense networks can rely upon a massive storage capacity, as $\alpha\propto K^{P-1}/N$, the price to pay is the need for large volumes of examples for training as, e.g. in the worst scenario of unsupervised learning, $M_{c}\propto 1/r^{2P}$ unsup . On the contrary, in the supervised setting, the presence of a teacher allows the standard regime $M_{c}\propto 1/r^{2}$ Lucibello .